# Decentralized Convergence to Equilibrium Prices in Trading Networks

### v1

@@ Abstract
We propose a decentralized market model in which agents can negotiate bilateral contracts. This builds on a similar, but centralized, model of trading networks introduced by Hatfield et al. (2013). Prior work has established that fully-substitutable preferences guarantee the existence of competitive equilibria which can be centrally computed. Our motivation comes from the fact that prices in markets such as over-the-counter markets and used car markets arise from decentralized negotiation among agents, which has left open an important question as to whether equilibrium prices can emerge from agent-to-agent bilateral negotiations. We design a best response dynamic intended to capture such negotiations between market participants. We assume fully substitutable preferences for market participants. In this setting, we provide proofs of convergence for sparse markets (covering many real world markets of interest), and experimental results for more general cases, demonstrating that prices indeed reach equilibrium, quickly, via bilateral negotiations. Our best response dynamic, and its convergence behavior, forms an important first step in understanding how decentralized markets reach, and retain, equilibrium.
@@ 1 Introduction
Section 4 complements our theoretical results with multi-agent simulations to develop a better quantitative understanding of convergence in our market dynamic. Specifically, we demonstrate the convergence, and speed of convergence, in markets with tree topologies, and move beyond the theoretical results to also demonstrate convergence in general (graph-based) markets. We explore numerous market configurations, including various market sizes and agent compositions, to analyze convergence and equilibrium properties. Our experiments demonstrate that equilibrium is reached much faster than the identified bounds. The experiments also highlight natural and desirable properties of our market dynamic, such as a general increase in welfare and satisfaction as we reach equilibrium, and the increase in utility as competition reduces (and vice versa). We also show how shocks propagate through the market, with knock-on effects before converging to a new equilibria, with the convergence speed governed by the severity of the shock.
@@ 2.1 Preferences
As in Hatfield et al. (2013), agents have heterogeneous and private preferences. An agent’s private valuation function $v^{i}$ maps every possible bundle $\Psi\subseteq\Omega_{i}$ of her trades to an integer value, and the empty bundle $\emptyset$ to 00.222Values for certain bundles can be negative, e.g. when they represent production costs for a seller or intermediary agent, as illustrated in Example 1. The utility $u^{i}(\Psi,\bm{p})$ an agent has for bundle $\Psi$ at prices $\bm{p}\in\mathbb{Z}^{\Omega_{i}}$ is the agent’s value for $\Psi$ plus the income from selling trades $\Psi_{i\rightarrow}$ minus the spending on buying trades $\Psi_{i\leftarrow}$: [^3] Heterogeneous preferences and quasilinear utilities are standard in the literature on substitutes markets and auctions (Hatfield et al. 2013; Kelso and Crawford 1982; Ausubel 2006).
@@
The roaster from Fig. 1 interested in sourcing raw beans from any one of the two producers, but not both, can express her valuation $v$ as $v(\emptyset)=0$, $v(\{\omega_{1}\})=v(\{\omega_{2}\})>0$ and $v(\{\omega_{1},\omega_{2}\})=-M$, where $\omega_{1}$ and $\omega_{2}$ denote the buying trades of the roaster and $M$ is a sufficiently large negative number. Moreover, the roaster may have the technological constraint that she can only participate in a selling trade (e.g. ground coffee) if she also participates in the buying trade (e.g. coffee beans). She can implement this by assigning value $-M$ to any technologically infeasible bundle of trades, e.g., $v(\{\rho\})=-M$, where $\rho$ denotes any trade where the roaster acts as a seller.
@@ 3 The Emergence of Equilibria
In extensive numerical experiments, we observe that FSMs always converge to equilibrium. We now present theoretical results towards a proof that FSMs converge almost surely. [^6] Recall that the randomness in the market arises from the choice of unsatisfied agent in each round of the market dynamic. This qualitative result is complemented by our experiments in Section 4, which focus primarily on showing the convergence speed and additional welfare analysis. All formal statements and proofs can be found in the extended online version (Lock et al. 2024). Let $V$ be an upper bound on the absolute value of the agents’ valuation functions and initial offers.
@@
We prove the single-trade case by contradiction: suppose that the dynamic does not terminate. Then we observe that one agent always accepts her counterpart’s offer for the trade $\omega$, while the other always rejects. So the two agent’s offers both increase, or both decrease, by $\epsilon$ after each round of best responses from both agents. But the buyer (seller) demands $\emptyset$ instead of the trade if the price is high (low) enough, a contradiction. In the two-trade case, we generalize this approach and use full substitutability to reach the required contradiction. We refer to the online appendix (Lock et al. 2024) for the full proof of Proposition 5.
@@
In the first step, we describe a procedure to construct a best response sequence $S=S^{1}S^{2}\ldots$ consisting of subsequences $S^{k}$. Each $S^{k}$ contains only agents from $J^{k}$, and all agents in $J^{k}$ are satisfied after applying sequence $S^{1}\ldots S^{k}$. Our second step is then to prove that $S$ is finite and the market terminates after applying $S$. We now sketch out each step in more detail; the full proof is given in the online appendix.
@@ 4 Experiments
We analyse several FSM topologies, ranging from bipartite networks of buyers and sellers to general networks with buyers, sellers, and intermediaries.
@@
Examples of these topologies, and their construction processes, are described and visualized in the extended version (Lock et al. 2024). The BS topology models direct trade between buyers and sellers, e.g., direct-to-consumer or business-to-consumer markets. The other topologies, which include intermediaries, can model a number of more complex markets, such as used car markets with buyers and sellers interacting through dealers (Hatfield et al. 2013), energy markets with consumers and generators mediated through suppliers (Morstyn, Teytelboym, and McCulloch 2018), or even general supply chains (Ostrovsky 2008).
@@ 4.1 Experimental Results
Convergence paths are visualized in Fig. 4(a). Increasing the number of agents increases the time to convergence, as there is more room for bargaining/trade among the agents (e.g. a buyer has access to more sellers, so has more negotiation opportunity). Importantly, in each case, we see convergence (the satisfied rate converges to $1$). Convergence rates across $N$ are visualized in Fig. 9(a).
@@
Figure 4: Convergence paths by market topology. The $x$-axis shows the number of iterations (e.g. number of best response updates), and the $y$-axis shows the proportion of satisfied agents.
@@
Introducing intermediaries into the network alters the convergence rates as transactions must happen through an intermediary. We explore different numbers of intermediaries, again, with a fixed connectivity rate of $r=0.1$, where for the possible pairs $i,j$ at least one of $i$ or $j$ must be an intermediary (e.g. there is no direct connection between buyers and sellers).
@@
Convergence rates across $\lambda$ are visualized in Fig. 4(c). Despite featuring cycles, these trading networks still converge, at a rate approximately linear with $\lambda$ (Fig. 9(c)). While we do not yet have theoretical guarantees of convergence under this setting, these results help to show that experimentally these general topologies also converge, opening a promising line of future research building on Theorem 10. The welfare of the agents does not change drastically across $\lambda$, but generally, as $\lambda$ increases the proportion of intermediaries in the network increases, so the welfare for buyers/sellers increases, and the welfare for intermediaries decreases (see extended version).
@@ 4.2 Shocks
Figure 6: The impact of exogenous shocks in a BS network. Fig. 6(a) shows the propagation of shocks measured as the additional proportion of unsatisfied agents. Fig. 6(b) shows the normalised reconvergence speed (as a proportion of the original convergence speed). Each $\circ$ represents the mean impact for a shock size (darker colour = larger magnitude). The filled region shows the effect range across these shock sizes.
To understand how exogenous shocks impact the market, shocks are applied to an agent’s value $c_{i}$, resampling to a new value in $C$. Shocks have a specific size $s$, which controls the variation from the original valuation, i.e., a shock size of $0.1$ means the updated $c_{i,t}$ will be in the range $[c_{i,t-1}*0.9,c_{i,t-1}*1.1]$ (clipped to the values in $C$). Shocks are applied to different proportions of agents in the network to analyse their propagation. Shocks only occur after the dynamics have converged to understand how these perturbations affect an otherwise converged system.
The results for the BS network are visualized in Fig. 6. We look at two key quantities: the proportion of additional (non-shocked) agents affected (Fig. 6(a)), e.g. due to knock-on effects of one borrower updating their valuation, and the normalized time taken to reconverge (Fig. 6(b)).
Propagation. When considering the propagation of shocks (Fig. 6(a)), a knock-on effect is observed in all cases, with additional non-shocked agents becoming affected (i.e. becoming unsatisfied and needing to update their prices). For example, when applying a shock to $25\%$ of the agents, approximately $50\%+$ of the agents become impacted and must update their responses. Generally speaking, the larger the shock magnitude, the more propagation throughout the network is observed, agreeing with real world insights on shock propagation’s, for example in production networks (Huneeus 2018) and global trade-investments (Starnini, Boguñá, and Serrano 2019), where the magnitude of the shock determines its impact on the overall system.
Reconvergence. The additional time to converge (Fig. 6(b)) increases linearly with the proportion of shocked agents and the shock magnitude, showing the disruption effects. As expected by the theory, irrespective of the shock size, the dynamics eventually settle down and reconverge to a new equilibrium. This reconvergence is significantly faster than the original convergence speed (Fig. 6(b)), essentially serving as a warm start to the convergence process, which is particularly true for small shock magnitudes.
@@ 4.3 Results Summary
These experimental resultssupport and extend the theoreticalguarantees in Section 3, by simulating the market dynamic for various trading networks and establishing consistent convergence across market topologies. The best response sequences constructed in the proof of Proposition 9 can be exponentially long. By contrast, we show experimentally that these markets converge rapidly. Additionally, demonstrate that more complex network structures also still converge under the proposed market dynamic,pening a promising line of future analysis extending Proposition 9. The experimental analysis provided insights into the resulting welfare and shock impacts across various market topologies, expanding the understanding of such market processes.
@@ A.2 Markets with more than two agents
These valuations implicitly endow each agent $i\in J$ with the opportunity to purchase (any subset of) the trades in $\Omega^{ext}_{i}$ at ‘frozen prices’ ${\sigma}^{-i}_{\omega}$, their counterparts’ offers, if this maximizes their utility; even though these trades do not appear in the market itself. For internal agents, $\widehat{v}^{i}=v^{i}$ as $\Omega^{ext}_{i}=\emptyset$. We note that the bundle $\Psi$ maximizing (4) is unique because of the way we perturb valuations to break ties. Moreover, Hatfield et al. (2015, Theorem 6) show that $\widehat{v}^{i}$ is fully- as long as $v^{i}$ is. We also map the state $(U,\bm{\sigma})$ of the original market to state $(\widehat{U},\widehat{\bm{\sigma}})$ of the restricted market by letting $\widetilde{U}=U\cap\Omega^{int}$ and $\widehat{\bm{\sigma}}$ be the offers $\bm{\sigma}$ restricted to the trades $\Omega^{int}$.
@@
This definition captures the notion of a merger of $J^{k}$, as the agents in this set can conduct internal trades for free (as the cost of buyer and the income of the seller cancel out). So the set of agents $J^{k}$ chooses a set $\Psi\subseteq\Omega^{k}$ of ‘interior’ trades that maximizes its aggregate valuation function. See also Hatfield et al. (2015, Section 5.2), where the valuations of merged agents is first introduced. In particular, Hatfield et al. (2015, Theorem 7) show that merging multiple fully- agents results in a fully- agent, so agents $1$ and $2$ in the two-agent market are both fully-substitutable.
@@ B.1 Network Construction
We consider Erdős-Rényi (ER) networks $G(N,\lambda)$, where $N$ is the number of traders, and $r=\frac{\lambda}{N}$ is the probability of each possible connection being made between the traders (i.e. $\lambda=N$ means that the graph is fully connected, and $\lambda\approx 1$ means that the graph is sparsely connected and tree-like).
@@
Trades are constructed both ways between two intermediaries (e.g. they are both interested in being the buyer or the seller).

### v2

@@ Abstract
We propose a decentralized market model in which agents can negotiate bilateral contracts. This builds on a similar, but centralized, model of trading networks introduced by Hatfield et al. in 2013. Prior work has established that fully-substitutable preferences guarantee the existence of competitive equilibria which can be centrally computed. Our motivation comes from the fact that prices in markets such as over-the-counter markets and used car markets arise from decentralized negotiation among agents, which has left open an important question as to whether equilibrium prices can emerge from agent-to-agent bilateral negotiations. We design a best response dynamic intended to capture such negotiations between market participants. We assume fully substitutable preferences for market participants. In this setting, we provide proofs of convergence for sparse markets (covering many real world markets of interest), and experimental results for more general cases, demonstrating that prices indeed reach equilibrium, quickly, via bilateral negotiations. Our best response dynamic, and its convergence behavior, forms an important first step in understanding how decentralized markets reach, and retain, equilibrium.
@@ 1 Introduction
Section 4 complements our theoretical results with multi-agent simulations to develop a better quantitative understanding of convergence in our market dynamic. [^2] The code for our simulations is available upon request from the authors. Specifically, we demonstrate the convergence, and speed of convergence, in markets with tree topologies, and move beyond the theoretical results to also demonstrate convergence in general (graph-based) markets. We explore numerous market configurations, including various market sizes and agent compositions, to analyze convergence and equilibrium properties. Our experiments demonstrate that equilibrium is reached much faster than the identified bounds. The experiments also highlight natural and desirable properties of our market dynamic, such as a general increase in welfare and satisfaction as we reach equilibrium, and the increase in utility as competition reduces (and vice versa). We also show how shocks propagate through the market, with knock-on effects before converging to a new equilibria, with the convergence speed governed by the severity of the shock.
@@ 2.1 Preferences
As in Hatfield et al. (2013), agents have heterogeneous and private preferences. An agent’s private valuation function $v^{i}$ maps every possible bundle $\Psi\subseteq\Omega_{i}$ of her trades to an integer value, and the empty bundle $\emptyset$ to 00.333Values for certain bundles can be negative, e.g. when they represent production costs for a seller or intermediary agent, as illustrated in Example 1. The utility $u^{i}(\Psi,\bm{p})$ an agent has for bundle $\Psi$ at prices $\bm{p}\in\mathbb{Z}^{\Omega_{i}}$ is the agent’s value for $\Psi$ plus the income from selling trades $\Psi_{i\rightarrow}$ minus the spending on buying trades $\Psi_{i\leftarrow}$: [^4] Heterogeneous preferences and quasilinear utilities are standard in the literature on substitutes markets and auctions (Hatfield et al. 2013; Kelso and Crawford 1982; Ausubel 2006).
@@
The roaster from Fig. 1 interested in sourcing raw beans from any one of the two producers, but not both, can express her valuation $v$ as $v(\emptyset)=0$, $v(\{\omega_{1}\})=v(\{\omega_{2}\})>0$ and $v(\{\omega_{1},\omega_{2}\})=-M$, where $\omega_{1}$ and $\omega_{2}$ denote the buying trades of the roaster and $M$ is a sufficiently large negative number. Moreover, the roaster may have the technological constraint that she can only participate in a selling trade (e.g., ground coffee) if she also participates in the buying trade (e.g., coffee beans). She can implement this by assigning value $-M$ to any technologically infeasible bundle of trades, e.g., $v(\{\rho\})=-M$, where $\rho$ denotes any trade where the roaster acts as a seller.
@@ 3 The Emergence of Equilibria
In extensive numerical experiments, we observe that FSMs always converge to equilibrium. We now present theoretical results towards a proof that FSMs converge almost surely. [^7] Recall that the randomness in the market arises from the choice of unsatisfied agent in each round of the market dynamic. This qualitative result is complemented by our experiments in Section 4, which focus primarily on showing the convergence speed and additional welfare analysis. All formal statements and proofs can be found in the appendix. Let $V$ be an upper bound on the absolute value of the agents’ valuation functions and initial offers.
@@
We prove the single-trade case by contradiction: suppose that the dynamic does not terminate. Then we observe that one agent always accepts her counterpart’s offer for the trade $\omega$, while the other always rejects. So the two agent’s offers both increase, or both decrease, by $\epsilon$ after each round of best responses from both agents. But the buyer (seller) demands $\emptyset$ instead of the trade if the price is high (low) enough, a contradiction. In the two-trade case, we generalize this approach and use full substitutability to reach the required contradiction. We refer to the appendix for the full proof of Proposition 5.
@@
In the first step, we describe a procedure to construct a best response sequence $S=S^{1}S^{2}\ldots$ consisting of subsequences $S^{k}$. Each $S^{k}$ contains only agents from $J^{k}$, and all agents in $J^{k}$ are satisfied after applying sequence $S^{1}\ldots S^{k}$. Our second step is then to prove that $S$ is finite and the market terminates after applying $S$. We now sketch out each step in more detail; the full proof is given in the appendix.
@@ 4 Experiments
We analyze several FSM topologies, ranging from bipartite networks of buyers and sellers to general networks with buyers, sellers, and intermediaries.
@@
Examples of these topologies, and their construction processes, are described and visualized in the appendix. The BS topology models direct trade between buyers and sellers, e.g., direct-to-consumer or business-to-consumer markets. The other topologies, which include intermediaries, can model a number of more complex markets, such as used car markets with buyers and sellers interacting through dealers (Hatfield et al. 2013), energy markets with consumers and generators mediated through suppliers (Morstyn, Teytelboym, and McCulloch 2018), or even general supply chains (Ostrovsky 2008).
@@ 4.1 Experimental Results
Convergence paths are visualized in Fig. 4(a). Increasing the number of agents increases the time to convergence, as there is more room for bargaining/trade among the agents (e.g., a buyer has access to more sellers, so has more negotiation opportunity). Importantly, in each case, we see convergence (the satisfied rate converges to $1$). Convergence rates across $N$ are visualized in Fig. 9(a).
@@
Figure 4: Convergence paths by market topology. The $x$-axis shows the number of iterations (e.g., number of best response updates), and the $y$-axis shows the proportion of satisfied agents.
@@
Introducing intermediaries into the network alters the convergence rates as transactions must happen through an intermediary. We explore different numbers of intermediaries, again, with a fixed connectivity rate of $r=0.1$, where for the possible pairs $i,j$ at least one of $i$ or $j$ must be an intermediary (e.g., there is no direct connection between buyers and sellers).
@@
Convergence rates across $\lambda$ are visualized in Fig. 4(c). Despite featuring cycles, these trading networks still converge, at a rate approximately linear with $\lambda$ (Fig. 9(c)). While we do not yet have theoretical guarantees of convergence under this setting, these results help to show that experimentally these general topologies also converge, opening a promising line of future research building on Theorem 10. The welfare of the agents does not change drastically across $\lambda$, but generally, as $\lambda$ increases the proportion of intermediaries in the network increases, so the welfare for buyers/sellers increases, and the welfare for intermediaries decreases (see appendix).
@@ 4.2 Shocks
Figure 6: The impact of exogenous shocks in a BS network. Fig. 6(a) shows the propagation of shocks measured as the additional proportion of unsatisfied agents. Fig. 6(b) shows the normalized reconvergence speed (as a proportion of the original convergence speed). Each $\circ$ represents the mean impact for a shock size (darker color = larger magnitude). The filled region shows the effect range across these shock sizes.
To understand how exogenous shocks impact the market, shocks are applied to an agent’s value $c_{i}$, resampling to a new value in $C$. Shocks have a specific size $s$, which controls the variation from the original valuation, i.e., a shock size of $0.1$ means the updated $c_{i,t}$ will be in the range $[c_{i,t-1}*0.9,c_{i,t-1}*1.1]$ (clipped to the values in $C$). Shocks are applied to different proportions of agents in the network to analyze their propagation. Shocks only occur after the dynamics have converged to understand how these perturbations affect an otherwise converged system.
The results for the BS network are visualized in Fig. 6. We look at two key quantities: the proportion of additional (non-shocked) agents affected (Fig. 6(a)), e.g., due to knock-on effects of one borrower updating their valuation, and the normalized time taken to reconverge (Fig. 6(b)).
Propagation. When considering the propagation of shocks (Fig. 6(a)), a knock-on effect is observed in all cases, with additional non-shocked agents becoming affected (i.e.,becoming unsatisfied and needing to update their prices). For example, when applying a shock to $25\%$ of the agents, approximately $50\%+$ of the agents become impacted and must update their responses. Generally speaking, the larger the shock magnitude, the more propagation throughout the network is observed, agreeing with real world insights on shock propagation’s, for example in production networks (Huneeus 2018) and global trade-investments (Starnini, Boguñá, and Serrano 2019), where the magnitude of the shock determines its impact on the overall system.
Reconvergence. The additional time to converge (Fig. 6(b)) increases approximately linearly with the proportion of shocked agents and the shock magnitude, showing the disruption effects. As expected by the theory, irrespective of the shock size, the dynamics eventually settle down and reconverge to a new equilibrium. This reconvergence is significantly faster than the original convergence speed (Fig. 6(b)), essentially serving as a warm start to the convergence process, which is particularly true for small shock magnitudes.
@@ 4.3 Results Summary
These experimental results support and extend the theoretical guarantees in Section 3, by simulating the market dynamic for various trading networks and establishing consistent convergence across market topologies. The best response sequences constructed in the proof of Proposition 9 can be exponentially long. By contrast, we show experimentally that these markets converge rapidly. Additionally, we demonstrate that more complex network structures also still converge under the proposed market dynamic, opening a promising line of future analysis extending Proposition 9. The experimental analysis provided insights into the resulting welfare and shock impacts across various market topologies, expanding the understanding of such market processes.
@@ A.2 Markets with more than two agents
These valuations implicitly endow each agent $i\in J$ with the opportunity to purchase (any subset of) the trades in $\Omega^{ext}_{i}$ at ‘frozen prices’ ${\sigma}^{-i}_{\omega}$, their counterparts’ offers, if this maximizes their utility; even though these trades do not appear in the market itself. For internal agents, $\widehat{v}^{i}=v^{i}$ as $\Omega^{ext}_{i}=\emptyset$. We note that the bundle $\Psi$ maximizing (4) is unique because of the way we perturb valuations to break ties. Moreover, Hatfield et al. (2015, Theorem 6) show that $\widehat{v}^{i}$ is fully-substitutable as long as $v^{i}$ is. We also map the state $(U,\bm{\sigma})$ of the original market to state $(\widehat{U},\widehat{\bm{\sigma}})$ of the restricted market by letting $\widetilde{U}=U\cap\Omega^{int}$ and $\widehat{\bm{\sigma}}$ be the offers $\bm{\sigma}$ restricted to the trades $\Omega^{int}$.
@@
This definition captures the notion of a merger of $J^{k}$, as the agents in this set can conduct internal trades for free (as the cost of buyer and the income of the seller cancel out). So the set of agents $J^{k}$ chooses a set $\Psi\subseteq\Omega^{k}$ of ‘interior’ trades that maximizes its aggregate valuation function. See also Hatfield et al. (2015, Section 5.2), where the valuations of merged agents is first introduced. In particular, Hatfield et al. (2015, Theorem 7) show that merging multiple fully-substitutable agents results in a fully-substitutable agent, so agents $1$ and $2$ in the two-agent market are both fully-substitutable.
@@ B.1 Network Construction
We consider Erdős-Rényi (ER) networks $G(N,\lambda)$, where $N$ is the number of traders, and $r=\frac{\lambda}{N}$ is the probability of each possible connection being made between the traders (i.e.,$\lambda=N$ means that the graph is fully connected, and $\lambda\approx 1$ means that the graph is sparsely connected and tree-like).
@@
Trades are constructed both ways between two intermediaries (e.g., they are both interested in being the buyer or the seller).